I am having trouble understanding how to find 2 real #'s between -2pi and 2pi that determine the point on the unit circle. i know 1 time around the unit circle is 2pi and one time around the unit circle clockwise is -2pi. I also need help to determine suppement of 5pi/12 and complement of 3pi/8. I am looking at my book and teacher's notes and I don't see it explained at all. Thanks so kindly for your time.
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(57397)

Find 2 real #'s between -2pi and 2pi that determine the point on the unit circle.
---
You have give us some information about the point.
=============================================================
I know 1 time around the unit circle is 2pi and one time around the unit circle clockwise is -2pi.
===========================
I also need help to determine supplement of
5pi/12
---
Ans: 2pi - (5/12)pi = (24/12)pi-(5/12)pi = (19/12)pi
=======================================================
The complement of 3pi/8
---
Ans: (1/2)pi - (3/8)pi
= (4/8)pi - (3/8)pi
---
= (1/8)pi
==============
Cheers,
Stan H.

Answer by Alan3354(30993)

I also need help to determine suppement of 5pi/12 and complement of 3pi/8
------
Supplement of 5pi/12 = 12pi/12 - 5pi/12 = 7pi/12
---------
Complement of 3pi/8 = 4pi/8 - 3pi/8 = pi/8